one number is greater than another number x by 5. also the sum of thrice the larger and twice the other number is 50. find the numbers
Answer by jean14(9)

Let x =smaller no.
X+5= larger no.
thrice the larger; that is 3(x+5)
Twice another number = 2x
We make equation. If the sum of 2numbers is 50 then ,we write
3(x+5) + 2x = 50
3x+15 +2x = 50 solve for x
5x= 50-15
5x/5= 35/5
X= 7 is the smaller no.
solve for larger number by substituting;
X+5= 7+5= 12 is the large no.
Checking
Subs. To equation 3(x+5) + 2x =50
3(7+5) + 2(7)= 50
3(12) +14 =50
36+14=50
50 =50 true
The numbers are 7 and 12.